OpenStudy (anonymous):

x/a-1=y/b solve for x

OpenStudy (lynfran):

x/(a-1)=y/b multiply both side by (a-1) we get x=(y/b)*(a-1)

OpenStudy (anonymous):

thank you so much

OpenStudy (lynfran):

welcome